OpenStudy (anonymous):

Simplify. Show your work. 5 1/3 +(-3 9/18)

OpenStudy (shinalcantara):

\[5\frac{ 1 }{ 3 } - 3 \frac{ 9 }{ 18 }\] change it first into improper fraction for easier solution

OpenStudy (shinalcantara):

after changing into improper fraction, get its LCD

OpenStudy (shinalcantara):

what did you get?

OpenStudy (anonymous):

8/1 - 21/9 @shinalcantara ?

OpenStudy (shinalcantara):

try to recheck.. in finding the improper fraction, you'll have: 

OpenStudy (anonymous):

63/18 @shinalcantara

OpenStudy (shinalcantara):

yep

OpenStudy (shinalcantara):

\[\frac{ 16 }{ 3 } - \frac{ 63 }{ 18 }\] now find the LCD

OpenStudy (anonymous):

3

OpenStudy (shinalcantara):

what number that when divided by both 3 and 18 will have a whole number as a quotient?

OpenStudy (shinalcantara):

is 3/18 be having a whole number as a quotient?

OpenStudy (shinalcantara):

wouldn't it be 18?

OpenStudy (shinalcantara):

\[\frac{ 16 }{ 3 }-\frac{ 63 }{ 18 } =\frac{ 6(16)-63 }{ 18 }\]

OpenStudy (shinalcantara):

18 divided by 3 makes 6. 6 is multiplied to the numerator which is 16. 18 divided by 18 is 1 multiplied to 63 is still 63. now you simplify

OpenStudy (anonymous):

32/21 @shinalcantara

OpenStudy (shinalcantara):

recheck.. what's 6*16?

OpenStudy (anonymous):

its 96 @shinalcantara

OpenStudy (shinalcantara):

\[\frac{ 96-63 }{ 18 }\] now subtract for you to have a value for the numerator

OpenStudy (anonymous):

96-63=33 @shinalcantara

OpenStudy (shinalcantara):

that makes the answer\[\frac{ 33 }{ 18 } \] now reduce this to lowest term by dividing both 33 and 18 by 3